the label the store brand mouthwash says that it is 22.3% alcohol by volume. the label on the popular store brand says that it is 12% alcohol by volume. if macy mixes 8 ounces of the store brand mouthwash with 5 ounces of the popular brand name mouthwash, what is the percent alcohol of the mixture?
Answer by stanbon(75887)

the label the store brand mouthwash says that it is 22.3% alcohol by volume. the label on the popular store brand says that it is 12% alcohol by volume. if macy mixes 8 ounces of the store brand mouthwash with 5 ounces of the popular brand name mouthwash, what is the percent alcohol of the mixture?
----
Equation:
alch + alch = alch
0.223*8 + 0.12*5 = x(8+5)
-----
223*8 + 120*5 = x(13000)
---------------------------
x = 0.1834 = 18.34%
=====================
Cheers,
Stan H.
================